# Quantum State Tomography of Photonic Qubits with Realistic Coherent Light Sources

### v1

@@ I Introduction
Quantum state tomography (QST) is a fundamental tool in quantum information science, providing a means to fully characterize the quantum state of a physical system [1, 2, 3]. It serves as a bridge between experimental measurements and the theoretical description of quantum systems, enabling us to verify and validate the behavior of quantum devices. Accurate knowledge of quantum states is essential for a wide range of applications, including quantum computing, quantum communication, and quantum cryptography [5, 6, 7, 8, 9]. As these technologies advance, the ability to reliably reconstruct quantum states from experimental data becomes increasingly important.
However, the practical implementation of QST is challenged by various imperfections that arise during the transmission of quantum states. Although photonic states are preferred for their practical applications [10, 11, 12], they encounter limitations when transmitted through optical fibers. As quantum states propagate through optical channels, they are subject to attenuation, which reduces the number of photons that can reach the detection system, and to various types of noise, such as shot noise and Raman scattering [14, 13, 15]. These factors can significantly distort the measured data, leading to inaccuracies in the reconstructed quantum states. Attenuation in optical fibers, for instance, causes an exponential decay in the signal strength with distance, while shot noise introduces random fluctuations in the photon count due to the inherent quantum nature of light. For this reason, developing and testing frameworks that account for various sources of measurement uncertainty remains an important area of ongoing research, see, for example, Refs. [16, 17, 18, 19].
@@ II Mathematical modeling of realistic sources
$m_{j}=\lceil\widetilde{\mathcal{N}}_{j}\>{\mathrm{Tr}}M_{j}\rho_{in}\rfloor,$ (3)
@@
$m_{j}=\lceil\widetilde{\mathcal{N}}_{j}\>{\mathrm{Tr}}M_{j}\rho_{in}+0.5\>(n_{j}^{R}+n_{j}^{cr})\rfloor,$ (8)

### v2

@@ I Introduction
Quantum state tomography (QST) is a fundamental tool in quantum information science, providing a means to fully characterize the quantum state of a physical system [1, 2, 3, 4]. It serves as a bridge between experimental measurements and the theoretical description of quantum systems, enabling us to verify and validate the behavior of quantum devices. Accurate knowledge of quantum states is essential for a wide range of applications, including quantum computing, quantum communication, and quantum cryptography [5, 6, 7, 8, 9, 10]. As these technologies advance, the ability to reliably reconstruct quantum states from experimental data becomes increasingly important.
However, the practical implementation of QST is challenged by various imperfections that arise during the transmission of quantum states. Although photonic states are preferred for their practical applications [11, 12, 13, 14], they encounter limitations when transmitted through optical fibers. As quantum states propagate through optical channels, they are subject to attenuation, which reduces the number of photons that can reach the detection system, and to various types of noise, such as shot noise and Raman scattering [16, 15, 17]. These factors can significantly distort the measured data, leading to inaccuracies in the reconstructed quantum states. Attenuation in optical fibers, for instance, causes an exponential decay in the signal strength with distance, while shot noise introduces random fluctuations in the photon count due to the inherent quantum nature of light. For this reason, developing and testing frameworks that account for various sources of measurement uncertainty remains an important area of ongoing research, see, for example, Refs. [18, 19, 20, 21].
@@ II Mathematical modeling of realistic sources
$m_{j}=\lceil\widetilde{\mathcal{N}}_{j}\>{\mathrm{Tr}}M_{j}\rho_{in}\rfloor\hskip 14.22636pt\text{for}\hskip 14.22636ptj=1,\dots,4,$ (3)
@@
$m_{j}=\lceil\widetilde{\mathcal{N}}_{j}\>{\mathrm{Tr}}M_{j}\rho_{in}+0.5\>(n_{j}^{R}+n_{j}^{cr})\rfloor\hskip 14.22636pt\text{for}\hskip 14.22636ptj=1,\dots,4,$ (8)